# Theoretical Insights into Fine-Tuning Attention Mechanism: Generalization and Optimization

### v1

@@ 3.1 Empirical Advantages of Fine-Tuning only Query,Value Matrices
Table 1 provides a detailed comparison of the impact of fine-tuning different weight matrices ($\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$) across various rank values $r$ and weight update strategies in LoRA fine-tuning on tasks like SST2, QNLI, QQP, and MNLI. As seen in the table, we can see a clear trend where solely updating the $\mathbf{W}_{v}$ matrix outperforms just learning the $\mathbf{W}_{q},\mathbf{W}_{k}$ matrix. Interestingly, the combination of fine-tuning both $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ often leads to performance that matches or even exceeds that achieved by fine-tuning all three matrices $\mathbf{W}_{q},\mathbf{W}_{k},$ and $\mathbf{W}_{v}$. This pattern is consistently observed across various tasks and rank values, further emphasizing the importance of these two matrices over $\mathbf{W}_{k}$ during fine-tuning.
Computational benefits. Here, we show that the reduced amount of adapted parameters by (roughly) $1/3$ provides computational gains. The key benefit of parameter-efficient method is to save memory during training, storage and communication [24]. Fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ alone as opposed to both $\mathbf{W}_{q}\&\mathbf{W}_{v}$ and $\mathbf{W}_{k}$ reduces the number of parameters by $1/3$, when the dimensions of $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ are the same.
@@ 3.2 Information-Theoretic Generalization Bounds
In the previous part, we establish that the Different Impact among $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ during fine-tuning. Some studies [30, 36, 3] often treat $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ as a single unit ($\mathbf{W}_{qk}=\mathbf{W}_{q}\mathbf{W}_{k}^{T}$), however, the benefits of fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ alone, rather than fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$, and $\mathbf{W}_{k}$ together, have yet to be further clarified. Therefore, we will further analyze this issue from an information-theoretic generalization perspective.
@@
Recalling A unified framework for parameter-efficient fine-tuning, we can model the fine-tuning process of the attention mechanism as $\mathbf{h}+\Delta\mathbf{h}=\mathbf{x}\mathbf{W}+\mathbf{x}\Delta\mathbf{W}$. Let $\mathbf{W}=\{\mathbf{W}_{i}\}_{i=1}^{L}$ be a set of abstract parameter matrices related to a pretrained model, where each $\mathbf{W}_{i}$ is associated with the parameters $\mathbf{W}_{q}^{i},\mathbf{W}_{k}^{i},\mathbf{W}_{v}^{i}$. The indices $1,...,L$ represent the layers of the model where these parameters are to be fine-tuned. Let $\mathcal{I}\subseteq\{1,...,L\}$ denote the subset of layers selected for fine-tuning. Given a fine-tuning training set $S_{N}$, let $r$ denote the chosen lora-rank, and assume each tuned parameter is quantized to $q$ bits. Define the following algorithmic frameworks for selecting an adaptation $\Delta\mathbf{W}=\{\Delta\mathbf{W}_{i}\}_{i=1}^{L}$ (with other details left open to choice).
(1) $\mathcal{A}_{QKV}$: For each $i\in\mathcal{I}$, optimize $\{\mathbf{W}_{q}^{i},\mathbf{W}_{k}^{i},\mathbf{W}_{v}^{i}\}_{i\in\mathcal{I}}$ to fit the data $S_{N}$.
(2) $\mathcal{A}_{QV}$: For each $i\in\mathcal{I}$, optimize $\{\mathbf{W}_{q}^{i},\mathbf{W}_{v}^{i}\}_{i\in\mathcal{I}}$ to fit the data $S_{N}$.
Then we have the following theorem to bound the generalization error using the information-theoretic generalization framework.
@@
Consider the algorithms of Definition 1. Assume the loss $\ell(\mathbf{W},Z)$ is R-subGaussian under $(\Delta\mathbf{W},{{Z}})\sim P_{\Delta\mathbf{W}\mid\mathbf{W}}\times\mu$. Then,
@@
where $\mathbf{W}_{q}^{i},\mathbf{W}_{k}^{i},\mathbf{W}_{v}^{i}\in\mathbb{R}^{d_{in}\times d_{out}}$. See Appendix B.2 for a proof.
@@ 4 Convergence Analysis in Optimization
4.1 An Insight into Inefficient Learning for Value Matrix
@@
Next, we present the second interesting phenomenon Efficient Convergence. We use the General Language Understanding Evaluation (GLUE, [40]) to evaluate the fine-tuning performance of different fine-tuning strategies, which consists of several language tasks that evaluate the understanding capabilities of language models. Using LoRA, we fine-tune Roberta-base from the RoBERTa family [26] and Llama3.1-8b [2] on MNLI, QQP, QNLI, and SST2 tasks with varying learning rates $(\eta_{QK},\eta_{V})$ to identify the optimal combination. Other empirical details are provided in Appendix C.1. We present our empirical results using LoRA to fine-tune language models, as visualized in the heatmaps (Figure 1 and Figure 2).
Figure 1: The test accuracy of RoBERTa-base fine-tuning was evaluated over 3 epochs for MNLI, QQP, and QNLI, and 6 epochs for SST-2, with a sequence length $T=128$ and using half-precision (FP16). The LoRA hyperparameters were set to $\alpha=r=8$. All reported values represent the average results across 3 random seeds. We highlight (1) the best overall accuracy and (2) the values where $\eta_{V}/\eta_{QK}=1$. These values are shown in red. For better visualization, when accuracy is lower than a fixed threshold, we set it to threshold.
In Figure 1 and Figure 2, we observe that (1) test accuracy consistently reaches its maximum for certain sets of learning rates where $\eta_{QK}<\eta_{V}$, outperforming the standard practice of setting $\eta_{QK}$ and $\eta_{V}$ equal. (2) More interestingly, the gap between the optimal choice of learning rates overall and the optimal choice when $\eta_{QK}=\eta_{V}$ varies across different tasks. This is probably due to the fact that harder task (like MNLI) requires more efficient feature learning. Additionally, we compare two optimal learning rate $(\eta_{QK},\eta_{V})$ settings in Figure 2 (Left), the $\eta_{V}>>\eta_{QK}$ setting has a better convergence than $\eta_{V}=\eta_{QK}$ setting in Figure 2 (Right).
It is also important to note that due to limited computational resources in our experiments, we use a sequence length of $T=128$ and fine-tune for only 3 epochs on MNLI and QQP. Therefore, it is expected that our test accuracies may be lower than those reported by Hu et al. [18], where the authors fine-tune RoBERTa-base with a sequence length of $T=512$ (for MNLI) and for more epochs (30 for MNLI). We do not include confidence intervals for clearer visualization, however, the fluctuations remain within acceptable limits. See Figure 2 (Right) for instance. In Appendix C.2, we provide additional results including the training loss.
@@ 4.2 Convergence Analysis for Learning Rate
It naturally raises the question of why $\eta_{QK}$ and $\eta_{V}$ should be set differently. In practice, most state-of-the-art models have a large width (embedding dimension), making it worthwhile to examine the training dynamics as the width approaches infinity.
@@
where $W^{*}\in\mathbb{R}^{n\times 1}$ are the fixed [^4] Here, we primarily focus on the case of $\Delta\mathbf{W}$ to provide insightful theoretical results. pre-trained weights, $b\in\mathbb{R},a\in\mathbb{R}^{1\times n}$ are adaptation weights, $x\in\mathbb{R}^{n}$ is the model input (This corresponds to $r=1$ in Definition 1). The training goal is to minimize the loss $\mathcal{L}(\theta)=\frac{1}{2}(f(x)-y)^{2}$ where $\theta=(a,b)$ and $(x,y)$ is an input-output datapoint [^5] To simplify the analysis, we assume that the fine-tuning dataset consists of a single sample, though our analysis can be easily generalized to multiple samples. All conclusions remain essentially valid when $(a,b)$ are matrices.. Similar to LoRA, we generally aim to initialize the product $a^{T}b$ to zero, ensuring that fine-tuning starts from the pre-trained model. This requires at least one of the weights, $a\ (\text{related to}\mathbf{W}_{q}\&\mathbf{W}_{k})$ or $b\ (\text{related to}\mathbf{W}_{v})$, to be initialized to zero. If both are initialized to zero, $\mathbf{W}_{q}\&\mathbf{W}_{k}$ learning cannot occur efficiently in init steps, as discussed in Section 4.1 (More detailed initialization settings are shown in Appendix B.3).
@@
$\Delta f_{t}$ $\triangleq f_{t}(x)-f_{t-1}(x)$
$=-\underbrace{\eta_{a}||x||^{2}U_{t-1}b_{t-1}^{2}}_{\delta_{t}^{1}}-\underbrace{\eta_{b}(xa_{t-1}^{T})^{2}U_{t-1}}_{\delta_{t}^{2}}+\underbrace{\eta_{a}\eta_{b}||x||^{2}(xa_{t-1}^{T})U_{t-1}^{2}b_{t-1}}_{\delta_{t}^{3}}.$
@@
Let us assume that we train the model with gradient descent with learning rate $\eta_{a}=\Theta(n^{c_{a}}),\eta_{b}=\Theta(n^{c_{b}})$ for some $c_{a},c_{b}\in\mathbb{R}$. In the study by Yang et al. [45], it is noted that the training dynamics primarily involve operations such as matrix-vector products and the summation of vectors or scalars. Given the nature of these operations, it is easy to see that any quantity in the training dynamics should be of order $n^{\gamma}$ for some $\gamma\in\mathbb{R}$. We write $v=\Theta(n^{\gamma[v]})$, for any quantity $v$ in the training dynamics. When $v$ is a vector, we use the same notation when all entries of $v$ are $\Theta(n^{\gamma[v]})$ (See Appendix B.4 for the formal definition of $\gamma$).
@@
In the case of Starting with a Toy setting, with $\eta_{a}=\Theta(n^{-1})$ and $\eta_{b}=\Theta(1)$, we have for all $t>1$, $i\in\{1,2,3\}$,$\delta_{t}^{i}=\Theta(1)$. In other words, the feature learning of attention mechanism is efficient when $\eta_{QK}(\eta_{a})=\Theta(n^{-1}),\eta_{V}(\eta_{b})=\Theta(1)$. We denote $\eta_{V}/\eta_{QK}$ as $\lambda$. We refer the reader to Appendix B.5 for more details on the proof.
@@
In practice, Theorem 2 implies that the learning rate for $\mathbf{W}_{v}$ should be generally much larger than that of $\mathbf{W}_{q}\&\mathbf{W}_{k}$ in fine-tuning. We verify that this scaling is valid for general neural network models in Section 4.1. Naturally, the optimal ratio $\lambda$ depends on the architecture and the fine-tuning task through the constants in ‘$\Theta$’. This represents a limitation of the asymptotic results, as they do not provide insights into how the task and neural architecture influence these constants. We will further address this issue in our future work.
5 An Example of Improving Fine-tuning Methods
Based on all our exciting insights, it becomes intuitive to design lightweight attention-based fine-tuning improvements, particularly for downstream tasks. To illustrate how theoretical analysis effectively guides experimental procedures, we propose an example method where we freeze the $\mathbf{W}_{k}$ and fine-tuning the $\mathbf{W}_{q}\&\mathbf{W}_{v}$ using different learning rates. This procedure is reported in Figure 5.
Experimental setup. We conduct experiments on widely adopted benchmark datasets [40] and Roberta-base model [26]. We selected two mainstream baselines: Full Fine-tuning and LoRA. Additionally, we adapt only the attention weights for downstream tasks, keeping the other modules frozen to maintain simplicity and validate the theoretical guidance through experiments. Moreover, as discussed in Remark 4, the optimal ratio $\lambda$ depends on the architecture and the fine-tuning task. In our experiments, we evaluated the performance for $\lambda$ values of 2, 4, and 8 (one can also determine a general optimal ratio through experiments, and even apply different settings across different layers of the model). We report the average results based on 3 random seeds, as shown in Table 2. The hyperparameter settings for the experiments can be found in Appendix C.1.2.
Method #Param (M) RTE STS-B MRPC CoLA MNLI SST-2 QQP QNLI
@@ 5 An Example of Improving Fine-tuning Methods
Results. We leverage our theoretical results (Theorem 1 and Theorem 2) to enhance the efficiency of existing fine-tuning methods, such as Full Fine-tune and LoRA, on downstream tasks. As shown in Table 2, the improved fine-tuning approach not only outperforms the original version but also significantly reduces the number of parameters. For instance, on the MRPC task, LoRA (QV) $r=16,\lambda=8$ (1.77M) achieves better performance compared to Full Fine-tune (QKV) (21.85M) and LoRA (QKV) $r=16$ (2.07M). This series of experiments clearly demonstrates that our theoretical insights effectively enhance fine-tuning algorithms, particularly in terms of memory usage and optimization efficiency. Moreover, these theoretical results can guide the improvement of other fine-tuning algorithms and even aid in the design of more efficient ones.

### v3

@@ 3.1 Empirical Advantages
Computational benefits. Here, we show that the reduced amount of adapted parameters by (roughly) $1/3$ provides computational gains. The key benefit of parameter-efficient method is to save memory during training, storage and communication [26]. Fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ alone as opposed to both $\mathbf{W}_{q}\&\mathbf{W}_{v}$ and $\mathbf{W}_{k}$ reduces the number of parameters by $1/3$, when the dimensions of $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ are the same. Moreover, we provide a discussion in Appendix B.2 about why fine-tune $\mathbf{W}_{q}\&\mathbf{W}_{v}$ instead of $\mathbf{W}_{k}\&\mathbf{W}_{v}$.
@@ 3.2 Information-Theoretic Generalization Bounds
In the previous part, we establish that the Unequal Importance of Attention Matrices among $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ during fine-tuning. Some studies [33, 4] often treat $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ as a single unit ($\mathbf{W}_{qk}=\mathbf{W}_{q}\mathbf{W}_{k}^{T}$), however, the benefits of fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ alone, rather than fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$, and $\mathbf{W}_{k}$ together, have yet to be further clarified. Therefore, we will further analyze this issue from an information-theoretic generalization perspective.
@@
Recalling A unified framework for parameter-efficient fine-tuning, we can model the fine-tuning process of the attention mechanism as $\mathbf{h}+\Delta\mathbf{h}=\mathbf{x}\mathbf{W}+\mathbf{x}\Delta\mathbf{W}$. Let $\mathbf{W}=\{\mathbf{W}_{i}\}_{i=1}^{L}$ be a set of abstract parameter matrices related to a pretrained model, where each $\mathbf{W}_{i}$ is associated with the parameters $\mathbf{W}_{q}^{i},\mathbf{W}_{k}^{i},\mathbf{W}_{v}^{i}$. The indices $1,...,L$ represent the layers of the model where these parameters are to be fine-tuned. Let $\mathcal{I}\subseteq\{1,...,L\}$ denote the subset of layers selected for fine-tuning. Given a fine-tuning training set $S_{N}$, let $r$ denote the chosen lora-rank, and assume each tuned parameter is quantized to $q$ bits. Define the following algorithmic frameworks for selecting an adaptation $\Delta\mathbf{W}=\{\Delta\mathbf{W}_{i}\}_{i=1}^{L}$ (with other details left open to choice). (1) $\mathcal{A}_{QKV}$: For each $i\in\mathcal{I}$, optimize $\{\mathbf{W}_{q}^{i},\mathbf{W}_{k}^{i},\mathbf{W}_{v}^{i}\}_{i\in\mathcal{I}}$ to fit the data $S_{N}$. (2) $\mathcal{A}_{QV}$: For each $i\in\mathcal{I}$, optimize $\{\mathbf{W}_{q}^{i},\mathbf{W}_{v}^{i}\}_{i\in\mathcal{I}}$ to fit the data $S_{N}$.
Then we use the information-theoretic generalization framework to bound the generalization error:
@@
Consider the algorithms of Definition 1. Assume the loss $\ell(\mathbf{W},Z)$ is R-subGaussian under $(\Delta\mathbf{W},{{Z}})\sim P_{\Delta\mathbf{W}\mid\mathbf{W}}\times\mu$. Then (See Appendix A.2 for a proof),
@@
where $\mathbf{W}_{q}^{i},\mathbf{W}_{k}^{i},\mathbf{W}_{v}^{i}\in\mathbb{R}^{d_{in}\times d_{out}}$.
@@ 4 Convergence Analysis in Optimization
4.1 An Insight into Inefficient Learning
@@
Next, we present the second interesting phenomenon Attention Matrices with Customized Learning Rate Lead to Better Convergence. We use the General Language Understanding Evaluation (GLUE, [44]) to evaluate the fine-tuning performance of different fine-tuning strategies, which consists of several language tasks that evaluate the understanding capabilities of language models. Using LoRA, we fine-tune Roberta-base from the RoBERTa family [28] and Llama3.1-8b [2] on MNLI, QQP, QNLI, and SST2 tasks with varying learning rates $(\eta_{QK},\eta_{V})$ to identify the optimal combination. Other empirical details are provided in Appendix B.1 and we evaluate the LLaMA3.1-8B model on more complex benchmarks in Appendix B.3.4. We present our empirical results using LoRA to fine-tune language models, as visualized in the heatmaps (Figure 1 and Figure 2).
Figure 1: The test accuracy of RoBERTa-base fine-tuning was evaluated over 3 epochs for MNLI, QQP, and QNLI, and 6 epochs for SST-2, with a sequence length $T=128$ and using half-precision (FP16). The LoRA hyperparameters were set to $\alpha=r=8$. All reported values represent the average results across 3 random seeds. We use red color to highlight (1) the best overall accuracy and (2) the values where $\eta_{V}/\eta_{QK}=1$. For better visualization, when accuracy is lower than a fixed threshold, we set it to threshold.
In Figure 1 and Figure 2 (Appendix B.3), we observe that (1) test accuracy consistently reaches its maximum for certain sets of learning rates where $\eta_{QK}<\eta_{V}$, outperforming the standard practice of setting $\eta_{QK}$ and $\eta_{V}$ equal. (2) More interestingly, the gap between the optimal choice of learning rates overall and the optimal choice when $\eta_{QK}=\eta_{V}$ varies across different tasks. This is probably due to the fact that harder task (like MNLI) requires more efficient feature learning. We compare two optimal learning rate $(\eta_{QK},\eta_{V})$ settings in Figure 2 (Left), the $\eta_{V}>>\eta_{QK}$ setting has a better convergence than $\eta_{V}=\eta_{QK}$ setting in Figure 2 (Right).
It is also important to note that due to limited computational resources in our experiments, we use a sequence length of $T=128$ and fine-tune for only 3 epochs on MNLI and QQP. Therefore, it is expected that our test accuracies may be lower than those reported by Hu et al. [20], where the authors fine-tune RoBERTa-base with a sequence length of $T=512$ (for MNLI) and for more epochs (30 for MNLI). We do not include confidence intervals for clearer visualization, however, the fluctuations remain within acceptable limits. See Figure 2 (Right) for instance. In Appendix B.3, we provide additional results including the training loss.
@@ 4.2 Convergence Analysis for Learning Rate
It naturally raises the question of why $\eta_{QK}$ and $\eta_{V}$ should be set differently. In practice, the large width (embedding dimension) of state-of-the-art models makes it valuable to examine training dynamics as width approaches infinity.
@@
where $W^{*}\in\mathbb{R}^{n\times 1}$ are the fixed [^6] Here, we primarily focus on the case of $\Delta\mathbf{W}$ to provide insightful theoretical results. pre-trained weights, $b\in\mathbb{R},a\in\mathbb{R}^{1\times n}$ are adaptation weights, $x\in\mathbb{R}^{n}$ is the model input (This corresponds to $r=1$ in Definition 1). The training goal is to minimize the loss $\mathcal{L}(\theta)=\frac{1}{2}(f(x)-y)^{2}$ where $\theta=(a,b)$ and $(x,y)$ is an input-output datapoint [^7] To simplify the analysis, we assume that the fine-tuning dataset consists of a single sample, though our analysis can be easily generalized to multiple samples. All conclusions remain essentially valid when $(a,b)$ are matrices.. Similar to LoRA, we generally aim to initialize the product $a^{T}b$ to zero, ensuring that fine-tuning starts from the pre-trained model. This requires at least one of the weights, $a\ (\text{related to}\mathbf{W}_{q}\&\mathbf{W}_{k})$ or $b\ (\text{related to}\mathbf{W}_{v})$, to be initialized to zero. If both are initialized to zero, $\mathbf{W}_{q}\&\mathbf{W}_{k}$ learning cannot occur efficiently in init steps, as discussed in Section 4.1 (More detailed initialization settings are in Appendix A.3).
@@
$\Delta f_{t}$ $\triangleq f_{t}(x)-f_{t-1}(x)=-\underbrace{\eta_{a}||x||^{2}U_{t-1}b_{t-1}^{2}}_{\delta_{t}^{1}}$
$-\underbrace{\eta_{b}(xa_{t-1}^{T})^{2}U_{t-1}}_{\delta_{t}^{2}}+\underbrace{\eta_{a}\eta_{b}||x||^{2}(xa_{t-1}^{T})U_{t-1}^{2}b_{t-1}}_{\delta_{t}^{3}}.$
@@
Let us assume that we train the model with gradient descent with learning rate $\eta_{a}=\Theta(n^{c_{a}}),\eta_{b}=\Theta(n^{c_{b}})$ for some $c_{a},c_{b}\in\mathbb{R}$. In the study by Yang et al. [47], it is noted that the training dynamics primarily involve operations such as matrix-vector products and the summation of vectors or scalars. Given the nature of these operations, it is easy to see that any quantity in the training dynamics should be of order $n^{\gamma}$ for some $\gamma\in\mathbb{R}$. We write $v=\Theta(n^{\gamma[v]})$, for any quantity $v$ in the training dynamics. When $v$ is a vector, we use the same notation when all entries of $v$ are $\Theta(n^{\gamma[v]})$ (See Appendix A.4 for the formal definition of $\gamma$).
@@
In the case of Starting with a Toy setting, with $\eta_{a}=\Theta(n^{-1})$ and $\eta_{b}=\Theta(1)$, we have for all $t>1$, $i\in\{1,2,3\}$,$\delta_{t}^{i}=\Theta(1)$. In other words, the feature learning of attention mechanism is efficient when $\eta_{QK}(\eta_{a})=\Theta(n^{-1}),\eta_{V}(\eta_{b})=\Theta(1)$. We denote $\eta_{V}/\eta_{QK}$ as $\lambda$. We refer the reader to Appendix A.5 for more details on the proof.
@@
In practice, Theorem 2 implies that the learning rate for $\mathbf{W}_{v}$ should be generally much larger than that of $\mathbf{W}_{q}\&\mathbf{W}_{k}$ in fine-tuning. We verify that this scaling is valid for general neural network models in Section 4.1. Naturally, the optimal ratio $\lambda$ depends on the architecture and the fine-tuning task through the constants in ‘$\Theta$’. This represents a limitation of the asymptotic results, as they do not provide insights into how the task and neural architecture influence these constants. We will further address this issue in future.
5 An Example of Improving Fine-tuning
Method Trainable #Param (M) RTE STS-B MRPC CoLA MNLI SST-2 QQP QNLI
Before Fine-tune 0 45.12 -3.18 66.66 1.09 32.95 49.31 44.72 50.81
@@ 5 An Example of Improving Fine-tuning
DoRA (QV) $r=8,\lambda=2$ 0.90 71.12 90.29 87.01 58.54 87.08 93.96 89.60 92.60
DoRA (QV) $r=8,\lambda=4$ 0.90 75.45 90.82 86.76 60.32 86.98 93.81 90.33 92.97
DoRA (QV) $r=8,\lambda=8$ 0.90 70.76 90.38 87.75 57.01 87.12 94.15 90.45 92.48
DoRA (QV) $r=16,\lambda=2$ 1.20 69.68 90.53 87.75 59.31 87.09 93.92 89.68 92.70
DoRA (QV) $r=16,\lambda=4$ 1.20 76.16 90.77 88.48 60.84 86.96 94.15 90.34 93.01
DoRA (QV) $r=16,\lambda=8$ 1.20 77.26 90.83 88.96 60.32 87.10 94.17 90.46 92.80
Based on all our exciting insights, it becomes intuitive to design lightweight attention-based fine-tuning improvements, particularly for downstream tasks. To illustrate how theoretical analysis effectively guides experimental procedures, we propose an example method where we freeze the $\mathbf{W}_{k}$ and fine-tuning the $\mathbf{W}_{q}\&\mathbf{W}_{v}$ using different learning rates. This procedure is reported in Figure 5. In Appendix B.2, we discuss how to set the ratio $\lambda$?
Experimental setup. We conduct experiments on widely adopted benchmark datasets [44] and Roberta-base model [28]. We selected mainstream baselines: Full Fine-tuning, LoRA [20] and DoRA [29]. Additionally, we adapt only the attention weights for downstream tasks, keeping the other modules frozen to maintain simplicity and validate the theoretical guidance through experiments. In our experiments, we evaluated the performance for $\lambda$ values of 2, 4, and 8 (one can also determine a general optimal ratio through experiments, and even apply different settings across different layers of the model). We report the average results based on 3 random seeds, as shown in Table 2. The hyperparameter settings for the experiments can be found in Appendix B.1.1 and the base model performance for each task can be seen in Table 2 and Appendix B.3.1. We also extend ablation experiments on different models (Mistral-7B [3]) in Appendix B.3.2.
Results. We leverage our theoretical results (Theorem 1 and Theorem 2) to enhance the efficiency of existing fine-tuning methods, such as Full Fine-tune, LoRA [20] and DoRA [29], on downstream tasks. As shown in Table 2, the improved fine-tuning approach not only outperforms the original version but also significantly reduces the number of parameters. For instance, on the MRPC task, LoRA (QV) $r=16,\lambda=8$ (1.77M) achieves better performance compared to Full Fine-tune (QKV) (21.85M) and LoRA (QKV) $r=16$ (2.07M). This series of experiments clearly demonstrates that our theoretical insights effectively enhance fine-tuning algorithms, particularly in terms of memory usage and optimization efficiency. Moreover, these theoretical results can guide the improvement of other fine-tuning algorithms and even aid in the design of more efficient ones.